A office supply store has been asked to determine weather or not the company should introduce a new computer desk next year. The store estimates that $6000 of new manufacturing equipment will need to be purchased and that the cost of constructing each desk will be $200. The store estimates that the revenue from each desk will be $450.
A) Determine the revenue function R (x) from the sale of x desks.
B) Determine the cost function C (x) for manufacturing x desks.
C) Find the break-even point.

Found 2 solutions by addingup, MathTherapy:
Answer by addingup(3677)

R = 450x-(6000+200x)
C = 6000+200x
BE = 450x = 6000+200x; 250x = 6000; x = 24 units

Answer by MathTherapy(10810)

A office supply store has been asked to determine weather or not the company should introduce a new computer desk next year. The store estimates that $6000 of new manufacturing equipment will need to be purchased and that the cost of constructing each desk will be $200. The store estimates that the revenue from each desk will be $450.
A) Determine the revenue function R (x) from the sale of x desks.
B) Determine the cost function C (x) for manufacturing x desks.
C) Find the break-even point.
R(x) = 450x
C(x) = 200x + 6,000
BEP equation: 450x = 200x + 6,000
Solve for x to get the BEP, in units.